OpenStudy (anonymous):

Factor 24p + 64 Answers. A. 24(p+40) B. 12(2p+5) C. 8(3p + 8) D. 3(8p + 21)

jimthompson5910 (jim_thompson5910):

what is the GCF of 24p and 64?

OpenStudy (anonymous):

8 right?

jimthompson5910 (jim_thompson5910):

Correct, so what do you get when you factor out 8?

OpenStudy (anonymous):

I'm not sure what you mean by factor out 8??

OpenStudy (anonymous):

Would it be 4??

jimthompson5910 (jim_thompson5910):

24 = 8*3 so 24p = 8*3p

jimthompson5910 (jim_thompson5910):

64 = 8*8

jimthompson5910 (jim_thompson5910):

that means 24p + 64 turns into 8*3p + 8*8 now you use the distributive property to factor out the GCF 8 to get 8(3p + 8)

jimthompson5910 (jim_thompson5910):

notice how if you were to distribute that 8 back through, you'd get 8(3p + 8) 8*3p + 8*8 24p + 64

OpenStudy (anonymous):

Oh okay:-DSo the answer would be C? Right

jimthompson5910 (jim_thompson5910):

Yes, the answer is C

OpenStudy (anonymous):

Okay thanks so much

jimthompson5910 (jim_thompson5910):

you're welcome